OpenStudy (anonymous):

solve the equation please 5t-5 ---- =70 -10

OpenStudy (anonymous):

t=????

OpenStudy (lgbasallote):

multiply -10 to both sides 5t - 5 = 70 x -10 5t - 5 = -700 can you solve for t now?

OpenStudy (lgbasallote):

sorry that's 5t+ 5 = 70 x -10 5t + 5 = -700 solve for t

OpenStudy (lgbasallote):

do you get that @Gerber96 ?

OpenStudy (anonymous):

I think so do I now try to find what multiplyed by 5t then + 5= -700????

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (anonymous):

5t(?)+5=-700

OpenStudy (lgbasallote):

no... 5t + 5 = -700 what should you do to get rid of that 5 on the left side?

OpenStudy (anonymous):

mulytiply 5 on both sides?

OpenStudy (lgbasallote):

if you multiply by 5 you will get 25t + 25 = -3500 you didnt make it simpler...try again

OpenStudy (anonymous):

divide?

OpenStudy (lgbasallote):

what will happen if you divide by 5?

OpenStudy (anonymous):

nvm got it thanks you definatley helped!!!!!

OpenStudy (lgbasallote):

<tips hat>